OpenStudy (anonymous):

At a grocery store, blueberries come packaged in 8-ounce containers for $2.80. At a farmer's market, blueberries cost $4.20 for 14 ounces. There are 16 ounces in one pound. Which statement accurately compares the cost of blueberries? Blueberries are $0.23 per pound more expensive at the farmer’s market. Blueberries are $0.50 per pound less expensive at the farmer’s market. Blueberries are $0.80 per pound less expensive at the farmer’s market. Blueberries are $1.40 per pound more expensive at the farmer’s market.

OpenStudy (texaschic101):

Grocery store.. 8 ounce containers for 2.80... 2.80/8 = 0.35 cents per ounce 16 ounces in a lb....so 16 * 0.35 = 5.60 per pound Farmers market .. 14 ounces for 4.20 4.20/14 = 0.30 cents per ounce 16 ounces in a lb....so 16 * 0.30 = 4.80 per pound 5.60 - 4.80 = 0.80 cents conclusion : farmers market is 0.80 cents less then the grocery store any questions ?

OpenStudy (anonymous):

thanks ! i suck at this stuff lol (: